The Cummings family wants to build a fence around their yard. If their yard is 182 m by 234 m, how many meters of fencing will they need?

Answer by fractalier(6550)

Perimeter of a rectangle P = 2L + 2W = 2(234) + 2(182) = 468 + 364 = 832 m